OpenStudy (anonymous):

Find the equation of the line that passes through (-3, 2) with a slope of (-4/5). Write your answer in Slope-Intercept Form. Answer

OpenStudy (anonymous):

y -2 = (-4/5) ( x + 3) -> y = -4x/5 - 12/5 + 2 = -4x/5 -1/2

OpenStudy (anonymous):

-2/5 ?????????

OpenStudy (anonymous):

my answer choices are -4/5x+5 -4/5x-2 , -4/5x-2/5, y=-4x-2...

OpenStudy (anonymous):

-12/5 + 2 = ( -12 + 10)/ 5 = -2/5

OpenStudy (anonymous):

Oops, -2/5 is right! I'm trying to simply and it's wrong. Thanks for reminding :")

OpenStudy (anonymous):

:) thanks for helpingg (:

OpenStudy (anonymous):

Just keep posting, you'll sure be benefited greatly from this website!